OpenStudy (anonymous):

5x^2+4x=39,998 I am solving for x

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

there are no like terms.

OpenStudy (anonymous):

\[x=\frac{1}{5} \left(-2\pm 19 \sqrt{554}\right) \]

OpenStudy (anonymous):

How did you get that? @robtobey

OpenStudy (anonymous):

x=89.04137744914263,−89.84137744914264

OpenStudy (anonymous):

How did you get that? @Dubba2000

OpenStudy (anonymous):

Step 1: Subtract 39998 from both sides

OpenStudy (anonymous):

Step 2: Use quadratic formula with a=5, b=4, c=-39998

OpenStudy (anonymous):

x= −b±sqrt(b2−4ac) 2a x= −(4)±sqrt((4)2−4(5)(−39998)) 2(5) x= −4±sqrt(799976) 10 x=89.04137744914263,−89.84137744914264

OpenStudy (anonymous):

Wait. my equation started out like this (2x+1)+x^2=40,000 and i got to 5x^2+4x=39,998 so idk what to do from there

OpenStudy (anonymous):

i will just work the original question really quick then

OpenStudy (anonymous):

okay thank you @Dubba2000

OpenStudy (anonymous):

x=199 or x=−201

OpenStudy (anonymous):

How? Can you explain? @Dubba2000

OpenStudy (anonymous):

Simplify both sides of the equation

OpenStudy (anonymous):

Subtract 40000 from both sides

OpenStudy (anonymous):

Factor left side of equation

OpenStudy (anonymous):

Set factors equal to 0

OpenStudy (anonymous):

1. x2+2x+1=40000 2. x2+2x+1−40000=40000−40000 x2+2x−39999=0 3. (x−199)(x+201)=0 4. x−199=0 or x+201=0 x=199 or x=−201

OpenStudy (anonymous):

(2x+1)^2+x^2=40,000

OpenStudy (anonymous):

\[5 x^2+4 x-39998=0 \]Apply the binomial theorem to the LHS of the above equation.\[\frac{-b\pm \sqrt{b^2-4 \text{ac}}}{2 a} \]

OpenStudy (anonymous):

hmm okay